### v2

@@ 10 Solving other problems
In this short section, we prove (or reprove) results that are easy consequences of our main result. We focus on decision problems; in the case of positive instances, explicit constructions can be done using Theorem 9.1.
@@ 10.1 Crossing number problem
We can solve the crossing number problem in $2^{\text{poly}(k)}n^{2}$ time.
@@
The reduction is known [13, Introduction] but we include it here for completeness. Consider the topological space that is a disk minus $k$ open disks with disjoint closures. To the boundary of each of these $k$ smaller disks, attach two edges whose endpoints are interleaved along the boundary. This topological space is homeomorphic to a two-dimensional simplicial complex $\mathscr{C}$ with $O(k)$ simplices.
@@
Actually, using Theorem 1.2, we obtain an algorithm that in $(k+w)^{O(k+w)}n$ time solves the crossing number problem when restricted to graphs of branchwidth at most $w$.
@@ 10.2 Embedding extension problem
We provide an algorithm for the embedding extension problem on complexes with no 3-book and with no isolated edge:
@@
We can solve the embedding extension problem for complexes where each edge of the complex is incident to either one or two triangles in $2^{\text{poly}(c+k)}n^{2}$ time.
@@
Without loss of generality, we can assume that the input complex $\mathscr{C}$ has no isolated vertex (see the proof of Proposition 3.1).
Let $U=\{u_{1},\ldots,u_{k}\}$, and let $p_{1},\ldots,p_{k}$ be the distinct points of $\mathscr{C}$ such that $u_{i}$ must be embedded at location $p_{i}$.
We augment $\mathscr{C}$ into a larger complex $\mathscr{C}^{\prime}$ as follows. For each $i$, let $\mathscr{S}_{i}$ be the orientable surface of Euler genus $10c+2i$. We attach $\mathscr{S}_{i}$ to $\mathscr{C}$ by connecting an arbitrary point $q_{i}$ of $\mathscr{S}_{i}$ with the point $p_{i}$ of $\mathscr{C}$ with a new segment $s_{i}$.
We augment $G$ into a larger graph $G^{\prime}$ as follows. For each $i$, let $H_{i}$ be a 2-connected graph with $O(c+i)$ vertices and edges that is embeddable on $\mathscr{S}_{i}$ but not on any surface (orientable or not) surface of smaller Euler genus. (It suffices to build graphs embedded on the surface with the prescribed genus with triangular faces, such that any non-contractible cycle has length at least four, by a result of Thomassen [40], and this is easy; see Figure 10.1.) We attach $H_{i}$ to $G$ by connecting an arbitrary vertex $h_{i}$ of $H_{i}$ with vertex $u_{i}$ via a new edge $e_{i}$.
@@
We claim that the instance of the embedding extension problem is positive if and only if $G^{\prime}$ is embeddable in $\mathscr{C}^{\prime}$, which concludes. Clearly, if the embedding extension problem is positive, then $G^{\prime}$ is embeddable in $\mathscr{C}^{\prime}$: Indeed, we embed $G$ into $\mathscr{C}$ as given by the solution of the embedding extension problem, we embed each graph $H_{i}$ to $\mathscr{S}_{i}$ mapping $h_{i}$ to $q_{i}$, and we embed each edge $e_{i}$ to the segment $s_{i}$.
Conversely, assume that $G^{\prime}$ is embeddable in $\mathscr{C}$. By induction on $i=k,k-1,\ldots,1$, by definition of $H_{i}$, by Lemma 3.4 and because $\mathscr{C}$ has no 3-book, we have that $H_{i}$ is embedded entirely into $\mathscr{S}_{i}$. Moreover, each face of $H_{i}$ in $\mathscr{S}_{i}$ is homeomorphic to a disk, because otherwise $H_{i}$ would be embeddable on a surface of lower genus.
Assume first that $u_{i}$ lies in $\mathscr{C}$. Because $h_{i}$ lies in $\mathscr{S}_{i}$, edge $h_{i}u_{i}$ goes through $p_{i}$. Because $\mathscr{C}$ has no isolated edge, we can pull $u_{i}$ to $p_{i}$, putting the other edges incident with $u_{i}$ in a neighborhood of the original edge $h_{i}u_{i}$.
Assume now that $u_{i}$ lies in $\mathscr{S}_{i}\cup s_{i}\setminus\{p_{i}\}$. We consider the connected component $G_{i}$ of $G$ containing $u_{i}$. The part of $G_{i}$ inside $\mathscr{S}_{i}\cup s_{i}$ lies in the union of a face of $H_{i}$ and of segment $s_{i}$, and is thus planar. If any other part of $G$ intersects $\mathscr{S}_{i}\cup s_{i}$, it is also planar and can be pulled out of $\mathscr{S}_{i}\cup s_{i}$. Then, the part of $G_{i}$ inside $\mathscr{S}_{i}$ can be pushed to a neighborhood of $p_{i}$, with $u_{i}$ on $p_{i}$.
Finally, we have an embedding of $G^{\prime}$ in which each $u_{i}$ lies on $p_{i}$; its restriction to $G$ satisfies the desired properties. ∎
With a bit more work, it should be possible to modify the dynamic program and the irrelevant vertex method to develop a direct solution to the embedding extension problem even in the presence of isolated edges. However, handling 3-books looks substantially more complicated, and we leave it as an open problem.
@@ 10.3 Planarity number problem
We can solve the planarity number problem in $2^{\text{poly}(c)}n^{2}$ time.
@@
Acknowledgment
We would like to thank Arnaud de Mesmay and Dimitrios Thilikos for useful discussions, and the anonymous reviewers of a previous version for their detailed comments.

### v3

@@ 10 Solving other problems
In this short section, we prove (or reprove) results that are easy consequences of our main result. We focus on decision problems; in the case of positive instances, explicit constructions can be done using Theorem 9.1 without overhead in the runtime. As we describe in Section 11, the decision problems can also be solved in time linear in $n$, but with an unspecified dependency in the parameter, by combining our techniques with Mohar’s linear-time algorithm to embed graphs into surfaces [35] together with a recent algorithm for the irrelevant vertex method [23].
@@ 10.1 Crossing number problem
We can solve the crossing number problem in $O(2^{\text{poly}(k)}n^{2})$ time.
@@
The reduction is known [15, Introduction] but we include it here for completeness. Consider the topological space that is a disk minus $k$ open disks with disjoint closures. To the boundary of each of these $k$ smaller disks, attach two edges whose endpoints are interleaved along the boundary. This topological space is homeomorphic to a two-dimensional simplicial complex $\mathscr{C}$ with $O(k)$ simplices.
@@
Actually, using Theorem 1.2, we obtain an algorithm that in $(k+w)^{O(k+w)}n$ time solves the crossing number problem when restricted to graphs of branchwidth at most $w$, which is a new result; see Lokshtanov, Panolan, Saurabh, Sharma, Xue, and Zehavi [31].
@@ 10.2 Embedding extension problem
We provide an algorithm for the embedding extension problem on complexes without 3-book:
@@
We can solve the embedding extension problem for complexes without 3-book in $O(2^{\text{poly}(c+k)}n^{2})$ time.
@@
Without loss of generality, we can assume that the input complex $\mathscr{C}$ has no isolated vertex (see the proof of Proposition 3.1). Let $u_{1},\ldots,u_{k}$ be vertices of the input graph $G$, and $p_{1},\ldots,p_{k}$ be distinct points of the input complex $\mathscr{C}$, such that $u_{i}$ must be embedded at location $p_{i}$.
We augment $\mathscr{C}$ into a larger complex $\mathscr{C}^{\prime}$, and $G$ into a larger graph $G^{\prime}$, as follows. For each $i$, let $H_{i}$ be a 2-connected graph of minimum degree at least three with $O(c+i)$ vertices and edges that is embeddable on the orientable surface of genus $10c+2i$ but not on any surface (orientable or not) of smaller Euler genus. (It suffices to build graphs embedded on the surface with the prescribed genus with triangular faces, such that any non-contractible cycle has length at least four, by a result of Thomassen [49], and this is easy; see Figure 10.1.) For each $i$, let $h_{i}$ be an arbitrary but fixed vertex of $H_{i}$. We take a copy $\mathscr{C}_{i}$ of $H_{i}$ and attach its copy of $h_{i}$ to point $p_{i}$ of $\mathscr{C}$. Let $\mathscr{C}^{\prime}$ be the resulting space, which can be represented as a complex made of $O(k(c+k))$ simplices. Similarly, for each $i$, we take a copy $G_{i}$ of $H_{i}$ and attach it to $G$, via the copy of $h_{i}$, to vertex $u_{i}$ of $G$. Let $G^{\prime}$ be the resulting graph.
@@
We claim that the instance of the embedding extension problem is positive if and only if $G^{\prime}$ is embeddable in $\mathscr{C}^{\prime}$, which concludes. Clearly, if the instance of the embedding extension problem is positive, then $G^{\prime}$ is embeddable in $\mathscr{C}^{\prime}$: Indeed, we embed $G$ into $\mathscr{C}$ as given by the solution of the embedding extension problem, and each graph $G_{i}$ to $\mathscr{C}_{i}$.
Conversely, assume that $G^{\prime}$ is embeddable in $\mathscr{C}$. By induction on $i=k,k-1,\ldots,1$, since $H_{i}$ is 2-connected, $G_{i}$ must be embedded either in $\mathscr{C}_{i}$ or in $\mathscr{C}$, and it must actually be in $\mathscr{C}_{i}$ by Lemma 3.4 and because $\mathscr{C}$ has no 3-book, and finally must occupy exactly $\mathscr{C}_{i}$ because it has minimum degree at least three. Necessarily, $u_{i}$ is mapped to $p_{i}$, except perhaps in the very special case where $u_{i}$ has degree zero in $G$, but in that case one can clearly modify the embedding of $G_{i}$ on $\mathscr{C}_{i}$ to ensure that property. Finally, taking the restriction of the embedding to $G$ gives the desired embedding. ∎
Handling 3-books looks substantially more complicated, and we leave it as an open problem.
@@ 10.3 Planarity number problem
We can solve the planarity number problem in $O(2^{\text{poly}(k)}n^{2})$ time.
@@
11 Conclusion
We conclude with a discussion on possible improvements on the running time of our algorithm.
First, as mentioned in the introduction, by combining our techniques with an algorithm to embed graphs on surfaces by Mohar [35] and a very recent algorithm for the irrelevant vertex method by Golovach, Kolliopoulos, Stamoulis, and Thilikos [23], we are able to get an algorithm whose running time has a linear dependency in $n$ for our problem, at the cost of an increase in the parameter dependency:
Theorem 11.1.
One can solve the embeddability problem of graphs into 2-dimensional simplicial complexes in time $f(c)\cdot O(n)$ for some (unspecified) function $f$, where $c$ is the number of simplices of the input 2-complex and $n$ is the total number of vertices and edges of the input graph.
In turn, this implies algorithms that are linear in $n$ (with an unspecified dependence in the parameter) in all applications of Section 10. It is, however, less clear whether one can obtain a similar runtime for computing actual embeddings (Theorem 9.1).
Proof.
We show that Theorem 1.3 holds with a runtime of $f(c)\cdot O(n)$. The result immediately follows by an application of Theorem 1.2. So we consider a 2-complex $\mathscr{C}$ with $c$ simplices, and a graph $G$ with $n$ vertices and edges in total. As we show next, the main result by Golovach et al. [23] applies to our problem and implies that one can find, in time $f(c)\cdot O(n)$, a set $I$ of vertices of $G$ such that GIG-I has branchwidth $O(c^{5/2})$, and such that $G$ embeds on $\mathscr{C}$ if and only if GIG-I does. This implies the theorem.
The result by Golovach et al. requires an embedding of $G$ on a surface of bounded genus. For this purpose, using Mohar’s algorithm [35], we compute an embedding of $G$ on a surface $\bar{\mathscr{S}}$ of genus $g\leq 10c$. If such an embedding does not exist, then by Lemma 3.4, we know that $G$ does not embed on $\mathscr{C}$.
The result by Golovach et al. applies to problems that satisfy the so-called insulation property. We show that it indeed holds in our problem (with no roots). So we assume that $G$ contains a so-called $r$-railed nest for $r=11c+4$, namely, cycles $C_{1},\ldots,C_{r}$ and paths $P_{1},\ldots,P_{r}$, such that, in the embedding of $G$ on some surface $\bar{\mathscr{S}}^{\prime}$, the cycles $C_{i}$ bound disks that are nested in this order, $C_{1}$ being the innermost cycle, and such that for every $i,j$, the cycle $C_{i}$ and the path $P_{j}$ intersect in a (possibly trivial) path. We have to prove that if GJG-J embeds on $\mathscr{C}$, then so does $G$, where $J$ is the set of vertices inside $C_{1}$ (in the embedding of $G$ on $\bar{\mathscr{S}}^{\prime}$).
The proof is an easy variation on that of Proposition 8.3. By removing at most $2c$ cycles and paths, we obtain that there is a circular wall of height $9c+4$ and circumference one that, in $\mathscr{C}$, does not intersect any singular vertex or isolated edge. Moreover, if $C_{9c+4}$ is the outer cycle, then the connected component of $G-C_{9c+4}$ that contains $J$ and the rest of the circular wall is planar. The only difference in the rest of the proof is that $J$ is not necessarily a single vertex $v$, but this does not affect the validity of the arguments. ∎
Finally, a second direction would be to improve the $2^{\text{poly}(c)}$ dependency while keeping the quadratic dependency in $n$. Our proof yields a $2^{O(c^{18+\varepsilon})}$ dependency; $18+\varepsilon$ is twice the exponent in the best known bound for the excluded grid theorem [11]. In particular, any improvement in that bound would improve the degree of the polynomial in $c$ in our algorithm. However, obtaining a much lower degree seems out of reach with our current techniques.
Acknowledgments
We would like to thank Petr Hliněný, Arnaud de Mesmay, Giannos Stamoulis, and Dimitrios Thilikos for useful discussions, and the anonymous reviewers for their detailed comments.